Market surveys show that there are two types of consumers for frozen yogurt. The first type like frozen yogurt and have an inverse demand curve of P =5 − 21Q. The second type is crazy about frozen yogurt and has an inverse demand curve of P = 20 − Q. In the town of Smallville, there are only 2 consumers: one of them likes frozen yogurt and the other is crazy about frozen yogurt.


Using the individual demand curves above, derive the market demand for frozen yogurt in Smallville. Plot the market demand curves.


2. Suppose that the market supply for frozen yogurt in Smallville is given by QS = 2+P. Find the equilibrium price and quantity. How much does each consumer buy at the equilibrium price?


1
Expert's answer
2021-02-17T17:21:32-0500

1. For the first consumer:

"P = 5 -21Q \\

21Q = 5 -P \\

Q_1 = \frac{5-P}{21}"

For the second consumer:

"P = 20 -Q \\

Q_2 = 20 -P"

Market demand curve:

"Q = Q_1 +Q_2 \\

= \frac{5-P}{21} + 20 -P \\

Q_D = 20.24 -1.05P"




2. At equilibrium

"Q_D = Q_S \\

20.24 -1.05P = 21P \\

2.05P = 18.24 \\

P = 8.91 \\

Q = 10.91"

At the equilibrium price, the first consumer does not buy at all as there is a too high price for him. Only the second consumer buys all the amount Q = 10.91. Ideally at price 8.91 he would like to purchase Q = 20 – P = 11.09, but as only 10.91 units are supplied, he would purchase that.